OpenStudy (anonymous):

find the surface area-- Square based Pyramid, with a slant height of 8, and the length of one side of the base = 5

OpenStudy (anonymous):

That would be (slant height x base x (1/2) x 4) + (base x base) = 80 + 25 =105 sq units